OpenStudy (anonymous):

Five more than a number tripled is half the number. What is the number? A.6 B.–2 C.–1 D.1

OpenStudy (jdoe0001):

let's use a number "n" 5 more than n n + 5 the number is tripled 3n + 5 all that is equal to half the number 3n+5 = 1/2n solve for "n"

OpenStudy (anonymous):

Would you do the reciprocal?

OpenStudy (jdoe0001):

you could multiply by the reciprocal of 1/2, sure

OpenStudy (anonymous):

It would all cancel out. It would be one, right?

OpenStudy (jdoe0001):

it'd cancel out the denominator on the right-hand side

OpenStudy (anonymous):

So you'd be left with 3n+5=n?

OpenStudy (jdoe0001):

\(\bf 3n+5 = \frac{1}{2}n \implies 3n+5 = \frac{n}{2}\\ 2 \times (3n+5) = 2 \times \left(\frac{n}{2}\right)\)

OpenStudy (anonymous):